# Light Control of Triplet Pairing in Correlated Electrons with Mixed-Sign Interactions

### v1

@@ II.1 Light-Enhanced Spin-Triplet Pairing Correlations
$\Delta^{p_{x}}_{\mathbf{i},0}$ $=\frac{1}{\sqrt{2}}\left(c_{\mathbf{i},\uparrow}c_{\mathbf{i}+\hat{x},\downarrow}+c_{\mathbf{i},\downarrow}c_{\mathbf{i}+\hat{x},\uparrow}\right)\,,$ (1)
@@
To further elucidate the anisotropic response of laser-induced $p$-wave pairing correlations, we map the pairing correlations onto the Bloch sphere at four representative time points: $t=-8t_{h}^{-1}$, $-2t_{h}^{-1}$, 00 and $7t_{h}^{-1}$ [see Fig. 2d]. At equilibrium, pairing correlations are isotropic without any symmetry breaking. However, once a linearly polarized pump is applied, this symmetry is broken, causing the Bloch sphere to shift preferentially toward the $p_{x}$ direction while suppressing the $p_{y}$ component. This asymmetry becomes more pronounced with higher pump fluence, indicating a stronger laser-induced anisotropy. After the pump, the Bloch sphere undergoes continuous oscillations along the horizontal axis, yet $\Phi_{x}(t)$ consistently dominates over $\Phi_{y}(t)$. Notably, the center of the Bloch sphere does not cross the origin, as shown by the directional arrows in Fig. 2d, reflecting sustained $p_{x}$ channel dominance triggered by the pump. The first maximal imbalance occurs near $t\sim\sigma=2.5t_{h}^{-1}$, with slight variations depending on the pump amplitude. Since the system is closed and lacks dissipation, the undamped laser energy drives continuous Bloch sphere oscillations over extended timescales, ensuring the persistence of the $p_{x}$ dominance and enhancement of spin-triplet pairing.
@@ II.2 Floquet Engineering of Anisotropic Spin Interactions
The resonance condition for Floquet-induced FM exchange is not confined to the narrow frequency window denoted by the red-shaded regime in Fig. 4b, but extends across a broader range of pump frequencies. This broadening arises from two primary effects. First, the above simplified Floquet analysis, which is based on atomic models, requires adjustments to incorporate band dispersion and quantum fluctuations in many-body systems. Second, the intermediate-state energy $E_{\mathrm{int}}$, which determines $J_{x}^{\rm eff}$ in Eq. (4), is not a fixed quantity in the doped EHM but varies across lattice sites in a given Fock state, depending on the local occupation configurations. As illustrated in the upper panels of Fig. 4c, configurations with isolated spin pairs or uniform neighbor densities yield $E_{\rm int}=U-V$, corresponding to a strong resonance near $\Omega\sim U-V=9t_{h}$ (solid lines in Fig. 3b). However, non-uniform occupations introduce a distribution of $E_{\rm int}$ values spanning from $U-4V$ to $U+2V$. At an average filling of $\langle n\rangle=0.5$, this variation extends the resonance window from $U+0.5V=7.5t_{h}$ to $U-2.5V=10.5t_{h}$, as indicated by dashed lines in Fig. 4b. This broadened resonance range aligns with the observed light-enhanced $p_{x}$ pairing instabilities in Fig. 3, reflecting the role of Floquet engineering in tuning pairing correlations.
@@ IV.1 Extended Hubbard Model
where $c_{i\sigma}(c^{\dagger}_{i\sigma})$ annihilates (creates) an electron at site $i$ with spin $\sigma=\uparrow,\downarrow$ and $n_{i\sigma}=c^{\dagger}_{i\sigma}c_{i\sigma}(n_{i}=n_{i\uparrow}+n_{i\downarrow})$ represents the electron density operator. The model incorporates three key parameters: the nearest-neighbor hopping amplitude $t_{h}$, the on-site Coulomb repulsion $U$, and the nearest-neighbor interaction $V$. In this study, the on-site $U$ is always repulsive, accounting for the double-occupation energy, while the nearest-neighbor interaction $V$ is attractive, representing an effective phonon-mediated interaction. We adopt model parameters $U=8t_{h}$ and $V=-t_{h}$, based on prior experimental estimates for 1D cuprate chains [32, 36]. For $t_{h}=300-400\,$meV, a 400-nm laser resides in the single-photon resonance and an 800-nm laser resides in the two-photon resonance. Simulations are conducted on a 4$\times$4 cluster with periodic boundary conditions.
@@ IV.2 Peierls substitution and the Krylov-Subspace Method
[…]

### v2

@@ II.1 Light-Enhanced Spin-Triplet Pairing Correlations
[…]
@@
To further elucidate the anisotropic response of laser-induced $p$-wave pairing correlations, we map the pairing correlations onto the Bloch sphere at four representative time points: $t=-8t_{h}^{-1}$, $2t_{h}^{-1}$, 0 and $7t_{h}^{-1}$ [see Fig. 2d]. At equilibrium, pairing correlations are isotropic without any symmetry breaking. However, once a linearly polarized pump is applied, this symmetry is broken, causing the Bloch sphere to shift preferentially toward the $p_{x}$ direction while suppressing the $p_{y}$ component. This asymmetry becomes more pronounced with higher pump fluence, indicating a stronger laser-induced anisotropy. After the pump, the Bloch sphere undergoes continuous oscillations along the horizontal axis, yet $\Phi_{x}(t)$ consistently dominates over $\Phi_{y}(t)$. Notably, the center of the Bloch sphere does not cross the origin, as shown by the directional arrows in Fig. 2d, reflecting sustained $p_{x}$ channel dominance triggered by the pump. The first maximal imbalance occurs near $t\sim\sigma=2.5t_{h}^{-1}$, with slight variations depending on the pump amplitude. Since the system is closed and lacks dissipation, the undamped laser energy drives continuous Bloch sphere oscillations over extended timescales, ensuring the persistence of the $p_{x}$ dominance and enhancement of spin-triplet pairing.
@@ II.2 Floquet Engineering of Anisotropic Spin Interactions
The resonance condition for Floquet-induced FM exchange is not confined to the narrow frequency window denoted by the red-shaded regime in Fig. 4b, but extends across a broader range of pump frequencies. This broadening arises from two primary effects. First, the above simplified Floquet analysis, which is based on atomic models, requires adjustments to incorporate band dispersion and quantum fluctuations in many-body systems. Second, the intermediate-state energy $E_{\mathrm{int}}$, which determines $J_{x}^{\rm eff}$ in Eq. (4), is not a fixed quantity in the doped EHM but varies across lattice sites in a given Fock state, depending on the local occupation configurations. As illustrated in the upper panels of Fig. 4c, configurations with isolated spin pairs or uniform neighbor densities yield $E_{\rm int}=U-V$, corresponding to a strong resonance near $\Omega\sim U-V=9t_{h}$ (solid lines in Fig. 3b). However, non-uniform occupations introduce a distribution of $E_{\rm int}$ values spanning from U4VU-4V to $U+2V$. At an average filling of $\langle n\rangle=0.5$, this variation extends the resonance window from $U+0.5V=7.5t_{h}$ to $U-2.5V=10.5t_{h}$, as indicated by dashed lines in Fig. 4b. This broadened resonance range aligns with the observed light-enhanced $p_{x}$ pairing instabilities in Fig. 3, reflecting the role of Floquet engineering in tuning pairing correlations.
@@ IV.1 Extended Hubbard Model
where $c_{i\sigma}(c^{\dagger}_{i\sigma})$ annihilates (creates) an electron at site $i$ with spin $\sigma=\uparrow,\downarrow$ and $n_{i\sigma}=c^{\dagger}_{i\sigma}c_{i\sigma}(n_{i}=n_{i\uparrow}+n_{i\downarrow})$ represents the electron density operator. The model incorporates three key parameters: the nearest-neighbor hopping amplitude $t_{h}$, the on-site Coulomb repulsion $U$, and the nearest-neighbor interaction $V$. In this study, the on-site $U$ is always repulsive, accounting for the double-occupation energy, while the nearest-neighbor interaction $V$ is attractive, representing an effective phonon-mediated interaction. We adopt model parameters $U=8t_{h}$ and $V=-t_{h}$, based on prior experimental estimates for 1D cuprate chains [32, 36]. For $t_{h}=300-400\,$meV, a 400-nm laser resides in the single-photon resonance and an 800-nm laser resides in the two-photon resonance. Simulations are conducted on a $\times4$ cluster with periodic boundary conditions.
@@ IV.2 Peierls substitution and the Krylov-Subspace Method
To describe the light-driven dynamics, we incorporate the Peierls substitution $t_{h}c^{\dagger}_{i\sigma}c_{j\sigma}\xrightarrow[]{}t_{h}e^{i\int_{r_{i}}^{r_{j}}{\textbf{A}(t)d\textbf{r}}}c^{\dagger}_{i\sigma}c_{j\sigma}$, which serves as an effective approximation of the light-matter interaction within the second-quantization framework. The specific pump vector potential $\textbf{A}(t)$ is provided by Eqs. (3) and (5) of the main text. We employ the SLEPc library to solve the ground state  [67], which yields a three-fold degenerate ground state under the chosen parameters. The time evolution of the ensemble-averaged density matrix is computed using the Krylov subspace method [68].